OpenStudy (anonymous):

multiply (√2)(-2√2+9) = -4+9√2 is my answer pls check

OpenStudy (anonymous):

jim_thompson5910 can u check multiply (√2)(-2√2+9) = -4+9√2 is my answer pls check

OpenStudy (anonymous):

-4 + 9*sqrt(2) This is the case b/c -2*sqrt(2) * sqrt(2) cancels the radicals and gives you -2 * 2 = 4 + sqrt(2)*9 Great work. http://www.tutorsean.net

OpenStudy (anonymous):

so sean it is postive 4 instead of -4+9√

OpenStudy (anonymous):

or is it 4 + √(2) * 9 the ans.